# Tensor global symmetries and the Stueckelberg mechanism for tensor fields

### v1

@@ 1 Introduction
Higher-form gauge symmetries frequently appear in theories such as supergravity and string theory, and they are long known and extensively studied [1, 2, 3], also from the mathematical standpoint of higher gauge theory [4, 5]. Recently, these symmetries were revisited as global symmetries through the lens of Noether’s theorem [6]. In this framework, “generalized” global symmetries extend standard symmetries by replacing rigid scalar transformation parameters with higher-form parameters.
@@
Higher-form global symmetries have been applied in various areas, including the formulation of relativistic hydrodynamics [20], studies in holography [21], and the construction of new models for topological phases of matter [22]. The latter application directly relates to the breaking of higher-form global discrete symmetries. This raises the broader question of spontaneous symmetry breaking in higher-form symmetries [23]. Just like spontaneous breaking of standard symmetry leads to a scalar field becoming the Goldstone boson of the broken symmetry, spontaneous breaking of $p$-form shift symmetries results in $p$-form gauge fields becoming the Goldstone bosons of the corresponding broken symmetries [13].
@@
A natural extension of this discussion is the generalization of $p$-form symmetries to mixed tensor symmetries (sometimes referred to as $(p,q)$-form symmetries). A notable example of a mixed symmetry tensor field is the graviton, which can be described as a $(1,1)$-symmetric field. This leads to an intriguing question: can the graviton be realized as a Goldstone boson for a spontaneously broken shift symmetry with a symmetric parameter? Recent works that explore this question are [24, 25, 26], see also [27] for a 1-form symmetry in nonlinear gravity. In [24], an explanation is offered by considering the linearized theory of gravity, focusing on a subclass of general shift symmetries where the shift symmetry parameter is written as the divergence of a $(2,1)$ parameter. Here, the invariant charges of the theory are constructed from the Riemann curvature, which involves second derivatives of the fields—a departure from typical currents in QFT that involve first derivatives.
In [26], an alternative formulation of linearized gravity is proposed to accommodate a general shift symmetry for the spin-2 field. In this formulation, the electric current is conserved only on-shell, however, the dual magnetic current is unusual: it is a reducible current that is not conserved under the right derivative. Moreover, neither the electric nor magnetic currents are gauge-invariant, complicating the standard gauging process, which typically involves coupling currents to background fields. However, this lack of gauge invariance is reminiscent of similar issues in the parent action formulation of linearized general relativity, where gauge invariance is restored with the help of an antisymmetric spin connection of Lorentz symmetry [28]. This suggests that incorporating a $2$-form field alongside the symmetric spin-2 field is essential to restore gauge invariance. This idea can be fully realized by using a first-order formulation of linearized gravity with a reducible $(1,1)$-tensor field, as originally shown in [29] and further studied in [28], offering an off-shell description of the electric-magnetic duality of linearized gravity. More recently, this was also extended to incorporate the duality relations of [30], which include a $\theta$-term, off-shell [31] and to provide a practical framework for tracking anomalies under duality [32].
@@
We illustrate this in three cases, primarily through the massive Curtright theory [33], which describes a massive $(2,1)$ mixed tensor field (the other two cases start with a massive graviton and a massive $(2,2)$ field, respectively). In this context, gauge invariance is restored by introducing a symmetric spin-2 field, an antisymmetric Kalb-Ramond field, and a vector field. The massless Curtright field possesses gauge symmetries with both symmetric and antisymmetric parameters. Spontaneous breaking of the symmetry generated by the symmetric parameter results in the realization of the graviton as the Goldstone boson of the broken symmetry. This perspective is advantageous because the $(2,1)$ Noether currents of the global shift symmetries are not only conserved on-shell but also gauge-invariant. This allows us to build a gauge-invariant theory from the ground up by coupling a free theory of Stueckelberg fields to a background Curtright field via the gauge-invariant current. Including nonminimal couplings, we identify a conserved and gauge invariant magnetic current, whose coupling generates a ’t Hooft anomaly for the symmetric shift symmetry.
@@ 2 Count of degrees of freedom
$\sum_{i=1}^{N}|\omega_{p_{i},q_{i}}|_{0}=\sum_{j=1}^{M}(|\Omega_{p_{j},q_{j}}|-|\Omega_{p_{j},q_{j}}|_{0})\,,$ (2)
@@
$\textrm{dim}_{D}(p,q)=\left(\begin{array}[]{c}D\\ p\end{array}\right)\left(\begin{array}[]{c}D+1\\ q\end{array}\right)\left(1-\frac{q}{p+1}\right)\,.$ (3)
@@
$|\omega_{p,q}|=\textrm{dim}_{D-1}(p,q)-\textrm{dim}_{D-1}(p-1,q-1)\,,$ (4)
$|\omega_{p,q}|_{0}=\textrm{dim}_{D-2}(p,q)-\textrm{dim}_{D-2}(p-1,q-1)\,,$ (5)
@@
$|\omega_{p}|=\left(\begin{array}[]{c}D-1\\ p\end{array}\right)\,,\quad|\omega_{p}|_{0}=\left(\begin{array}[]{c}D-2\\ p\end{array}\right)\,.$ (10)
@@
$|\omega_{p}|_{0}=|\Omega_{p+1}|-|\Omega_{p+1}|_{0}\,,$ (11)
@@
$|\phi+a|_{0}=\begin{cases}\,|A+B|-|A+B|_{0}\\[4.0pt] \,|H|-|H|_{0}\end{cases}\,.$ (12)
@@ 3 Higher form shift symmetries
$S[\omega]$ $=$ $-\frac{1}{2}\int\mathrm{d}\omega\wedge\ast\,\mathrm{d}\omega=-\frac{1}{2(p+1)!}\int\mathrm{d}^{D}x\,F_{\mu_{1}\dots\mu_{p+1}}F^{\mu_{1}\dots\mu_{p+1}}\,,$ (16)
@@
$J_{\text{e}}=\mathrm{d}\omega\,,\quad J_{\text{m}}=\ast\,\mathrm{d}\omega\,.$ (17)
@@
$S[\omega,\Omega_{\text{e}},\Omega_{\text{m}}]=\int\bigg{(}-\frac{1}{2}(\mathrm{d}\omega-\Omega_{\text{e}})\wedge\ast(\mathrm{d}\omega-\Omega_{\text{e}})+\Omega_{\text{m}}\wedge\mathrm{d}\omega\bigg{)}\,.$ (18)
@@
$\delta_{\text{e}}S=\int\varepsilon_{\text{e}}\wedge\mathrm{d}\Omega_{\text{m}}\,,$ (19)
@@
$\mathcal{I}=\int\mathrm{d}\Omega_{\text{m}}\wedge\mathrm{d}\Omega_{\text{e}}\,,$ (20)
@@
$S[\omega,\Omega_{\text{e}}]=-\frac{1}{2}\int\bigg{(}(\mathrm{d}\omega-\Omega_{\text{e}})\wedge\ast(\mathrm{d}\omega-\Omega_{\text{e}})+\mathrm{d}\Omega_{\text{e}}\wedge\ast\,\mathrm{d}\Omega_{\text{e}}\bigg{)}\,,$ (21)
which is gauge invariant under the symmetry generated by the spacetime-dependent parameter $\varepsilon_{\text{e}}$. It is easily recognizable that (21) is a generalized Stueckelberg action. For $p=0$ it is the usual one for a massive 1-form, and fixing the gauge by setting $\omega$ (in this case, a scalar) to zero, namely taking the unitary gauge, results in the Proca action (1). This could be described starting directly from the generalization of the Proca action for a massive $(p+1)$-form $\Omega$ with unit mass $m=1$ and performing the field redefinition
$\Omega\to\Omega-\mathrm{d}\omega\,.$ (22)
The field strength remains unchanged, as it does not see the field $\omega$, and the action (1) is recovered. Note though that this approach does not reveal the richer structure of conserved currents and generalized global symmetries, it does not give any information on ’t Hooft anomalies, and it is not suitable for electric/magnetic duality, which is not a feature of the massive theory. On the other hand, it is a guiding formulation when it is not a priori clear how gauging works, as we will see below.
@@ 4.1 Stueckelberg action for the massive graviton
$\hskip 22.0pt=-\frac{1}{4}\int\mathrm{d}^{4}x\int_{\text{B}}\bigg{(}\mathrm{d}H\star\mathrm{d}H-H\star H\bigg{)}\,.$ (23)
In the last line, we expressed the action in a compact and suggestive form. It is based on introducing anticommuting coordinates that are integrated out via Berezin integration $\int_{\text{B}}$. The mathematical details are described in [35], inspired and closely related to the bi-form formalism of [36, 37] and [38, 39], which transfers differential form notation to tensors of mixed symmetry. Since we use this in an auxiliary way in this paper and always write component expressions at the end, we only briefly explain the basic definitions for what appears in the action (23) and other related actions.
@@
$\theta^{\mu}\theta^{\nu}=-\theta^{\nu}\theta^{\mu}\,,\quad\widetilde{\theta}^{\mu}\widetilde{\theta}^{\nu}=-\widetilde{\theta}^{\nu}\widetilde{\theta}^{\mu}\,,\quad\theta^{\mu}\widetilde{\theta}^{\nu}=\widetilde{\theta}^{\nu}\theta^{\mu}\,.$ (24)
@@
$\omega=\frac{1}{p!q!}\omega_{[\mu_{1}\dots\mu_{p}][\nu_{1}\dots\nu_{q}]}(x)\theta^{\mu_{1}}\dots\theta^{\mu_{p}}\widetilde{\theta}^{\nu_{1}}\dots\widetilde{\theta}^{\nu_{q}}\,.$ (25)
@@
$\star\,\omega=\frac{1}{(D-p-q)!}\eta^{D-p-q}\,\widetilde{\omega}\,,$ (26)
@@
$\int_{\text{B}}:=\int\mathrm{d}^{D}\theta^{\mu}\mathrm{d}^{D}\widetilde{\theta}^{\mu}\,.$ (27)
@@
$\star\,\mathrm{d}H=\eta_{\mu_{1}\nu_{1}}\partial_{\nu_{2}}H_{\mu_{2}\nu_{3}}\theta^{\mu_{1}}\theta^{\mu_{2}}\widetilde{\theta}^{\nu_{1}}\widetilde{\theta}^{\nu_{2}}\widetilde{\theta}^{\nu_{3}}\,.$ (28)
@@
$\delta H_{\mu\nu}=\partial_{\mu}\alpha_{\nu}+\partial_{\nu}\alpha_{\mu}\,,$ (29)
@@
$\delta H_{\mu\nu}=-2\partial_{\mu}\partial_{\nu}\varphi\,,$ (30)
@@
$\delta\phi=\varphi\quad\text{and}\quad\delta a_{\mu}=\alpha_{\mu}\,.$ (31)
@@
$H_{\mu\nu}\to\mathring{F}_{\mu\nu}:=H_{\mu\nu}-\partial_{\mu}a_{\nu}-\partial_{\nu}a_{\mu}+2\partial_{\mu}\partial_{\nu}\phi\,.$ (32)
@@
$\delta\mathring{F}_{\mu\nu}=0\,.$ (33)
In the spirit of the Stueckelberg mechanism, substituting this into the Fierz-Pauli action results in a gauge invariant action for the massive graviton at the expense of introducing the additional massless fields. Since substituting this into the expanded form of the action gives a rather complicated result, let us present it in the compact form introduced above. This only requires to write (32) as
$\mathring{F}=H-\mathrm{d}\widetilde{a}-\widetilde{\mathrm{d}}a+2\mathrm{d}\widetilde{\mathrm{d}}\phi\,.$ (34)
@@
$\hskip 110.0pt-\,(H-\mathrm{d}\widetilde{a}-\widetilde{\mathrm{d}}a+2\mathrm{d}\widetilde{\mathrm{d}}\phi)\star(H-\mathrm{d}\widetilde{a}-\widetilde{\mathrm{d}}a+2\mathrm{d}\widetilde{\mathrm{d}}\phi)\bigg{)}\,.\qquad$ (35)
@@
$S_{\text{FP}}^{\text{St}}=\int\mathrm{d}^{4}x\int_{\text{B}}\bigg{(}-\frac{1}{4}\big{(}\mathrm{d}H\star\mathrm{d}H-H\star H\big{)}+\frac{1}{2}\big{(}\mathrm{d}\widetilde{a}\star\mathrm{d}\widetilde{a}+2H\star\mathrm{d}\widetilde{\mathrm{d}}\phi-2H\star\mathrm{d}\widetilde{a}\big{)}\bigg{)}\,.\,$ (36)
The first parenthesis is the Fierz-Pauli action (23). After these manipulations, we may perform the Berezin integration and write the final form of the remaining terms in components up to total derivatives:
$S_{\text{FP}}^{\text{St}}-S_{\text{FP}}=\int\mathrm{d}^{4}x\bigg{(}\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\partial_{\mu}a_{\nu}H^{\mu\nu}+\,\partial^{\mu}a_{\mu}H^{\nu}{}_{\nu}-\Box\phi H^{\mu}{}_{\mu}+\partial_{\mu}\partial_{\nu}\phi H^{\mu\nu}\bigg{)}\,,$ (37)
@@
This result teaches us an important lesson. If we want to start from the massless fields $\phi$ and $a$ and couple them to symmetric tensor background fields, the corresponding currents can be read off (37). Moreover, the starting point has to be different than simply the sum of the canonical kinetic terms for $\phi$ and $a$; total derivatives will play a role in this. Before presenting the details of this procedure, let us comment on the coupling of $H$ with $\phi$ and the apparent absence of the kinetic term for the scalar in (37). As already explained in [34], this can be accounted for by a Weyl rescaling of the graviton,
$H_{\mu\nu}\to\widehat{H}_{\mu\nu}=H_{\mu\nu}+\eta_{\mu\nu}\phi\,.$ (38)
@@
$S_{\text{FP}}^{\text{St}}\supset\int\mathrm{d}^{4}x\,\phi\,\Box\phi\,.$ (39)
Essentially, this is due to the fact that the mixing terms of $\phi$ and $H_{\mu\nu}$ in (37) generate the term $\phi R$, with $R$ the linearized curvature scalar for $H_{\mu\nu}$. The role of the scalar field here is very different from that of the usual Goldstone scalar. In the zero-coupling limit it decouples, unlike the zero coupling limit in gauge theory where the Goldstone boson kinetic term survives [34]. We will call such fields “secondary”, as further explained in Section 4.3.

### v2

@@ 1 Introduction
Higher-form gauge symmetries frequently appear in theories such as supergravity and string theory, and they are long known and extensively studied [1, 2, 3, 4], also from the mathematical standpoint of higher gauge theory [5, 6]. Recently, these symmetries were revisited as global symmetries through the lens of Noether’s theorem [7]. In this framework, “generalized” global symmetries extend standard symmetries by replacing rigid scalar transformation parameters with higher-form parameters.
@@
Higher-form global symmetries have been applied in various areas, including the formulation of relativistic hydrodynamics [21], studies in holography [22], and the construction of new models for topological phases of matter [23]. The latter application directly relates to the breaking of higher-form global discrete symmetries. This raises the broader question of spontaneous symmetry breaking in higher-form symmetries [24, 25]. Just like spontaneous breaking of standard symmetry leads to a scalar field becoming the Goldstone boson of the broken symmetry, spontaneous breaking of $p$-form shift symmetries results in $p$-form gauge fields becoming the Goldstone bosons of the corresponding broken symmetries [14].
@@
A natural extension of this discussion is the generalization of $p$-form symmetries to mixed tensor symmetries (sometimes referred to as $(p,q)$-form symmetries). A notable example of a mixed symmetry tensor field is the graviton, which can be described as a $(1,1)$-symmetric field. This leads to an intriguing question: can the graviton be realized as a Goldstone boson for a spontaneously broken shift symmetry with a symmetric parameter? Recent works that explore this question are [26, 27, 28], see also [29] for a 1-form symmetry in nonlinear gravity. There is an entirely different approach [30, 31, 32] to describe graviton as a Goldstone boson based on the nonlinear realizations of the affine and conformal symmetries which are then spontaneously broken to the Poincare group. In [26], an explanation is offered by considering the linearized theory of gravity, focusing on a subclass of general shift symmetries where the shift symmetry parameter is written as the divergence of a $(2,1)$ parameter. Here, the invariant charges of the theory are constructed from the Riemann curvature, which involves second derivatives of the fields—a departure from typical currents in QFT that involve first derivatives.
In [28], an alternative formulation of linearized gravity is proposed to accommodate a general shift symmetry for the spin-2 field. In this formulation, the electric current is conserved only on-shell, however, the dual magnetic current is unusual: it is a reducible current that is not conserved under the right derivative. Moreover, neither the electric nor magnetic currents are gauge-invariant under linearized diffeomorphism, complicating the standard gauging process, which typically involves coupling currents to background fields. However, this lack of gauge invariance is reminiscent of similar issues in the parent action formulation of linearized general relativity, where gauge invariance is restored with the help of an antisymmetric spin connection of Lorentz symmetry [33]. This suggests that incorporating a $2$-form field alongside the symmetric spin-2 field is essential to restore gauge invariance. This idea can be fully realized by using a first-order formulation of linearized gravity with a reducible $(1,1)$-tensor field, as originally shown in [34] and further studied in [33], offering an off-shell description of the electric-magnetic duality of linearized gravity. More recently, this was also extended to incorporate the duality relations of [35], which include a $\theta$-term, off-shell [36] and to provide a practical framework for tracking anomalies under duality [37].
@@
We illustrate this in three cases, primarily through the massive Curtright theory [38], which describes a massive $(2,1)$ mixed tensor field (the other two cases start with a massive graviton and a massive $(2,2)$ field, respectively); see [39, 40] for earlier works on the massive mixed tensor fields and [41, 42] for the discussion of the massive dual spin-2 field. In this context, gauge invariance is restored by introducing a symmetric spin-2 field, an antisymmetric Kalb-Ramond field, and a vector field. The massless Curtright field possesses gauge symmetries with both symmetric and antisymmetric parameters. Spontaneous breaking of the symmetry generated by the symmetric parameter results in the realization of the graviton as the Goldstone boson of the broken symmetry. This perspective is advantageous because the $(2,1)$ Noether currents of the global shift symmetries are not only conserved on-shell but also gauge-invariant. This allows us to build a gauge-invariant theory from the ground up by coupling a free theory of Stueckelberg fields to a background Curtright field via the gauge-invariant current. Including nonminimal couplings, we identify a conserved and gauge invariant magnetic current, whose coupling generates a ’t Hooft anomaly for the symmetric shift symmetry.
@@ 2 Count of degrees of freedom
$\sum_{i=1}^{N}|\omega_{p_{i},q_{i}}|_{0}=\sum_{j=1}^{M}(|\Omega_{p_{j},q_{j}}|-|\Omega_{p_{j},q_{j}}|_{0})\,,$ (2.2)
@@
$\textrm{dim}_{D}(p,q)=\left(\begin{array}[]{c}D\\ p\end{array}\right)\left(\begin{array}[]{c}D+1\\ q\end{array}\right)\left(1-\frac{q}{p+1}\right)\,.$ (2.3)
@@
$|\omega_{p,q}|=\textrm{dim}_{D-1}(p,q)-\textrm{dim}_{D-1}(p-1,q-1)\,,$ (2.4)
$|\omega_{p,q}|_{0}=\textrm{dim}_{D-2}(p,q)-\textrm{dim}_{D-2}(p-1,q-1)\,,$ (2.5)
@@
$|\omega_{p}|=\left(\begin{array}[]{c}D-1\\ p\end{array}\right)\,,\quad|\omega_{p}|_{0}=\left(\begin{array}[]{c}D-2\\ p\end{array}\right)\,.$ (2.10)
@@
$|\omega_{p}|_{0}=|\Omega_{p+1}|-|\Omega_{p+1}|_{0}\,,$ (2.11)
@@
$|\phi+a|_{0}=\begin{cases}\,|A+B|-|A+B|_{0}\\[4.0pt] \,|H|-|H|_{0}\end{cases}\,.$ (2.12)
@@ 3 Higher form shift symmetries
$S[\omega]$ $=$ $-\frac{1}{2}\int\mathrm{d}\omega\wedge\ast\,\mathrm{d}\omega=-\frac{1}{2(p+1)!}\int\mathrm{d}^{D}x\,F_{\mu_{1}\dots\mu_{p+1}}F^{\mu_{1}\dots\mu_{p+1}}\,,$ (3.1)
@@
$J_{\text{e}}=\mathrm{d}\omega\,,\quad J_{\text{m}}=\ast\,\mathrm{d}\omega\,.$ (3.2)
@@
$S[\omega,\Omega_{\text{e}},\Omega_{\text{m}}]=\int\bigg{(}-\frac{1}{2}(\mathrm{d}\omega-\Omega_{\text{e}})\wedge\ast(\mathrm{d}\omega-\Omega_{\text{e}})+\Omega_{\text{m}}\wedge\mathrm{d}\omega\bigg{)}\,.$ (3.3)
@@
$\delta_{\text{e}}S=\int\varepsilon_{\text{e}}\wedge\mathrm{d}\Omega_{\text{m}}\,,$ (3.4)
@@
$\mathcal{I}=\int\mathrm{d}\Omega_{\text{m}}\wedge\mathrm{d}\Omega_{\text{e}}\,,$ (3.5)
@@
$S[\omega,\Omega_{\text{e}}]=-\frac{1}{2}\int\bigg{(}(\mathrm{d}\omega-\Omega_{\text{e}})\wedge\ast(\mathrm{d}\omega-\Omega_{\text{e}})+\mathrm{d}\Omega_{\text{e}}\wedge\ast\,\mathrm{d}\Omega_{\text{e}}\bigg{)}\,,$ (3.6)
which is gauge invariant under the symmetry generated by the spacetime-dependent parameter $\varepsilon_{\text{e}}$. It is easily recognizable that (3.6) is a generalized Stueckelberg action. For $p=0$ it is the usual one for a massive 1-form, and fixing the gauge by setting $\omega$ (in this case, a scalar) to zero, namely taking the unitary gauge, results in the Proca action (2.1). This could be described starting directly from the generalization of the Proca action for a massive $(p+1)$-form $\Omega$ with unit mass $m=1$ and performing the field redefinition
$\Omega\to\Omega-\mathrm{d}\omega\,.$ (3.7)
The field strength remains unchanged, as it does not see the field $\omega$, and the action (2.1) is recovered. Note though that this approach does not reveal the richer structure of conserved currents and generalized global symmetries, it does not give any information on ’t Hooft anomalies, and it is not suitable for electric/magnetic duality, which is not a feature of the massive theory. On the other hand, it is a guiding formulation when it is not a priori clear how gauging works, as we will see below.
@@ 4.1 Stueckelberg action for the massive graviton
$\hskip 22.0pt=-\frac{1}{4}\int\mathrm{d}^{4}x\int_{\text{B}}\bigg{(}\mathrm{d}H\star\mathrm{d}H-H\star H\bigg{)}\,.$ (4.1)
In the last line, we expressed the action in a compact and suggestive form. It is based on introducing anticommuting coordinates that are integrated out via Berezin integration $\int_{\text{B}}$. The mathematical details are described in [44], inspired and closely related to the bi-form formalism of [45, 46] and [47, 48], which transfers differential form notation to tensors of mixed symmetry. Since we use this in an auxiliary way in this paper and always write component expressions at the end, we only briefly explain the basic definitions for what appears in the action (4.1) and other related actions.
@@
$\theta^{\mu}\theta^{\nu}=-\theta^{\nu}\theta^{\mu}\,,\quad\widetilde{\theta}^{\mu}\widetilde{\theta}^{\nu}=-\widetilde{\theta}^{\nu}\widetilde{\theta}^{\mu}\,,\quad\theta^{\mu}\widetilde{\theta}^{\nu}=\widetilde{\theta}^{\nu}\theta^{\mu}\,.$ (4.2)
@@
$\omega=\frac{1}{p!q!}\omega_{[\mu_{1}\dots\mu_{p}][\nu_{1}\dots\nu_{q}]}(x)\theta^{\mu_{1}}\dots\theta^{\mu_{p}}\widetilde{\theta}^{\nu_{1}}\dots\widetilde{\theta}^{\nu_{q}}\,.$ (4.3)
@@
$\star\,\omega=\frac{1}{(D-p-q)!}\eta^{D-p-q}\,\widetilde{\omega}\,,$ (4.4)
@@
$\int_{\text{B}}:=\int\mathrm{d}^{D}\theta^{\mu}\mathrm{d}^{D}\widetilde{\theta}^{\mu}\,.$ (4.5)
@@
$\star\,\mathrm{d}H=\eta_{\mu_{1}\nu_{1}}\partial_{\nu_{2}}H_{\mu_{2}\nu_{3}}\theta^{\mu_{1}}\theta^{\mu_{2}}\widetilde{\theta}^{\nu_{1}}\widetilde{\theta}^{\nu_{2}}\widetilde{\theta}^{\nu_{3}}\,.$ (4.6)
@@
$\delta H_{\mu\nu}=\partial_{\mu}\alpha_{\nu}+\partial_{\nu}\alpha_{\mu}\,,$ (4.7)
@@
$\delta H_{\mu\nu}=-2\partial_{\mu}\partial_{\nu}\varphi\,,$ (4.8)
@@
$\delta\phi=\varphi\quad\text{and}\quad\delta a_{\mu}=\alpha_{\mu}\,.$ (4.9)
@@
$H_{\mu\nu}\to\mathring{F}_{\mu\nu}:=H_{\mu\nu}-\partial_{\mu}a_{\nu}-\partial_{\nu}a_{\mu}+2\partial_{\mu}\partial_{\nu}\phi\,.$ (4.10)
@@
$\delta\mathring{F}_{\mu\nu}=0\,.$ (4.11)
In the spirit of the Stueckelberg mechanism, substituting this into the Fierz-Pauli action results in a gauge invariant action for the massive graviton at the expense of introducing the additional massless fields. Since substituting this into the expanded form of the action gives a rather complicated result, let us present it in the compact form introduced above. This only requires to write (4.10) as
$\mathring{F}=H-\mathrm{d}\widetilde{a}-\widetilde{\mathrm{d}}a+2\mathrm{d}\widetilde{\mathrm{d}}\phi\,.$ (4.12)
@@
$\hskip 110.0pt-\,(H-\mathrm{d}\widetilde{a}-\widetilde{\mathrm{d}}a+2\mathrm{d}\widetilde{\mathrm{d}}\phi)\star(H-\mathrm{d}\widetilde{a}-\widetilde{\mathrm{d}}a+2\mathrm{d}\widetilde{\mathrm{d}}\phi)\bigg{)}\,.\qquad$ (4.13)
@@
$S_{\text{FP}}^{\text{St}}=\int\mathrm{d}^{4}x\int_{\text{B}}\bigg{(}-\frac{1}{4}\big{(}\mathrm{d}H\star\mathrm{d}H-H\star H\big{)}+\frac{1}{2}\big{(}\mathrm{d}\widetilde{a}\star\mathrm{d}\widetilde{a}+2H\star\mathrm{d}\widetilde{\mathrm{d}}\phi-2H\star\mathrm{d}\widetilde{a}\big{)}\bigg{)}.$ (4.14)
The first parenthesis is the Fierz-Pauli action (4.1). After these manipulations, we may perform the Berezin integration and write the final form of the remaining terms in components up to total derivatives:
$S_{\text{FP}}^{\text{St}}-S_{\text{FP}}=\int\mathrm{d}^{4}x\bigg{(}\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\partial_{\mu}a_{\nu}H^{\mu\nu}+\,\partial^{\mu}a_{\mu}H^{\nu}{}_{\nu}-\Box\phi H^{\mu}{}_{\mu}+\partial_{\mu}\partial_{\nu}\phi H^{\mu\nu}\bigg{)},$ (4.15)
@@
This result teaches us an important lesson. If we want to start from the massless fields $\phi$ and $a$ and couple them to symmetric tensor background fields, the corresponding currents can be read off (4.15). Moreover, the starting point has to be different than simply the sum of the canonical kinetic terms for $\phi$ and $a$; total derivatives will play a role in this. Before presenting the details of this procedure, let us comment on the coupling of $H$ with $\phi$ and the apparent absence of the kinetic term for the scalar in (4.15). As already explained in [43], this can be accounted for by a Weyl rescaling of the graviton,
$H_{\mu\nu}\to\widehat{H}_{\mu\nu}=H_{\mu\nu}+\eta_{\mu\nu}\phi\,.$ (4.16)
@@
$S_{\text{FP}}^{\text{St}}\supset\int\mathrm{d}^{4}x\,\phi\,\Box\phi\,.$ (4.17)
Essentially, this is due to the fact that the mixing terms of $\phi$ and $H_{\mu\nu}$ in (4.15) generate the term $\phi R$, with $R$ the linearized curvature scalar for $H_{\mu\nu}$. The role of the scalar field here is very different from that of the usual Goldstone scalar. In the zero-coupling limit it decouples, unlike the zero coupling limit in gauge theory where the Goldstone boson kinetic term survives [43]. We will call such fields “secondary”, as further explained in Section 4.3.